7. A ball is equally likely to be painted yellow or blue. Two such balls are placed in an urn. Given that one of the balls is yellow, whats the probability that both balls are yellow? [10 pts] A - bath or Yellow B – one is Yellow ÞCA) POR ) D(Alb) = PCA)
Answer #1

P( blue) = 1/2

P(yellow) = 1/2

p( both yellow | 1 ball is yellow) = P(both yellow) / P(one is yellow)

= 1/2 *1/2 / 1/2*1/2+ 1/2*1/2 =1/4 / 1/2 = 2/4 = 1/2 =0.5